OpenStudy (anonymous):

A system of equations is shown below: 6x - 2y = 3 (equation 1) 5x + 3y = 4 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof?

OpenStudy (anonymous):

May i help?

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

KK 1 sec :)

OpenStudy (anonymous):

Okie. Do you want the answer options?

OpenStudy (anonymous):

She should plug in the same number for both equations and see :) fan and medal?

OpenStudy (anonymous):

OH YES PLEASE

OpenStudy (anonymous):

please give me the options @<3bae<3

OpenStudy (anonymous):

Show that the solution to the system of equations 10x + y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 10x – 2y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 11x – y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 11x + y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations

OpenStudy (anonymous):

?

OpenStudy (anonymous):

The 2 in the middle are subtraction for the first equation.

OpenStudy (anonymous):

A!! :)

OpenStudy (anonymous):

Medal and fan? and @ me for more help @<3bae<3 :)